The name of the book is basic coastal engineering. 3rd edition

8. A pressure gage located 1 m off the bottom in water 10 m deep measures an average maximum pressure of 100 kN/m2 having an8. d-10m 1.0m Pmax = 100kN/m2 T-12s cosh kd Cos ( ot z 9.0m, d-2-10-9-1.0m cos (ka- ot) 1.0

8. A pressure gage located 1 m off the bottom in water 10 m deep measures an average maximum pressure of 100 kN/m2 having an average fluctuation period of 12 s. Determine the height and period of the wave causing the measured pressure fluctuation.
8. d-10m 1.0m Pmax = 100kN/m2 T-12s cosh kd Cos ( ot z 9.0m, d-2-10-9-1.0m cos (ka- ot) 1.0
Answer #1

Solution:
Given that

The following data should be considering.
Pressure gage located 1 m
The bottom in water 10 m deep
Avg maximum pressure of 100 KN/m^2
Fluctuation period of 12 s
p + dp = 100 kN/m2
=> dp = 100000 - (10*1000*9.8)
=> dp = 2000 N/m2

This dp corresponds to the height of the wave from the surface of water.

=> dp = 1000 * h * 9.8
=> h = 2000 / (1000*9.8)
=> h=2000/9800
=> h = 0.204 m
Thus the height of the wave = 0.204 m or 20.4 cm
the period will be twice as that of the avg. fluctuation period i.e., T = 2*12 = 24 s